solve the following equation for A : 2A/3=8+4A
Answer by stanbon(75887)

solve the following equation for A : 2A/3=8+4A
-------------------------------------
Multiply thru by 3 to get rid of the fraction.
2A = 24 + 12A
10A = -24
A = -12/5
---------------
Cheers,
Stan H.